Circular No. 8485

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET D/1819 W1 (BLANPAIN) AND 2003 WY_{25} Already more than a year ago, S. Foglia, Milan, Italy, reported a sug-

gestion by M. Micheli that backward integration of the orbit of 2003 WY_{25} given on MPEC 2003-Y78 (Catalina Sky Survey discovery announcement on MPEC 2003-W41) suggested possible identity — though showing discordances extending up to 17^{o} deg in ω — with the lost comet D/1819 W1 = 1819 IV, which was itself tentatively shown by H. B. Ridley (1957, BAA Circ. No. 382) to be related to the one-time Phoenicid meteor shower of 1956 Dec. 5. P. Jenniskens, NASA Ames Research Center, has now independently suggested the 1819–2003 identity with an ω discordance of 0^{o}.2. Computations by B. G. Marsden, Harvard-Smithsonian Center for Astrophysics, that included reexamination of the 1819–1820 observations confirm a best-fit gravitational linkage with ω discordance 0^{o}.2. He also showed that the discordances in all three angular elements can be reduced to 0^{o}.01 by starting from the following orbital elements for 2003 WY_{25} (which had H = 21.1 and was consistently of stellar appearance despite a passage only 0.025 AU from the earth on 2003 Dec. 12):

Epoch = 2003 Dec. 27.0 TT

T = 2003 Dec. 11.5776 TT  ω = 9.^{o}0695
e = 0.675583  Ω = 69.3827 }2000.0
q = 1.000069 AU  i = 5.9292
a = 3.082662 AU n^{o} = 0.1821022 P = 5.412 years

Although backward integration of this orbit gives T too late in 1819, adjustment by ΔT = −4.28 days and modification of the angular elements within the range indicated above yield the result

Epoch = 1819 Nov. 22.0 TT

T = 1819 Nov. 20.27 TT  ω = 349.^{o}65
e = 0.7028  Ω = 80.02 }2000.0
q = 0.8893 AU  i = 9.23
a = 2.9928 AU n^{o} = 0.19036 P = 5.18 years

which satisfactorily represents 10 of the 13 observations made at Paris, Bologna, and Milan during 1819 Dec. 14–1820 Jan. 15 within 90′′. The integrated orbital elements at the time of the Phoenicid shower are T = 1956 Oct. 25.32 TT, ω = 0^{o}.14, Ω = 74^{o}.37, i = 9^{o}.60 (equinox 2000.0), q = 0.9914 AU, e = 0.6767, a = 3.0669 AU, P = 5.37 years.

2005 February 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green